Jun throws a stone straight upward alongside a tree with an initial velocity of 32 ft/sec.
The stone rises until it is even with the top of the tree and then falls back to the ground. How tall is the
tree?
1
Expert's answer
2020-11-27T10:56:29-0500

Let`s start with recognising given valuables:

Vinitial=32ft/sec,

And we need to find:

hmax=?

We have given everything to calculate hmax through below equation:

hmax=Vinitial2/(2*g), g=9.81 m/sec2=32.18 ft/sec2.

hmax=322/(2*32.18)=15.91 ft.

Then the answer is hmax=15.91 feet.

The length of the Tree is equal to hmax.